# Coarsened confounding for causal effects: a large-sample framework

### v1

@@ Abstract
There has been widespread use of causal inference methods for the rigorous analysis of observational studies and to identify policy evaluations. In this article, we consider coarsened exact matching, developed in Iacus et al. (2011). While they developed some statistical properties, in this article, we study the approach using asymptotics based on a superpopulation inferential framework. This methodology is generalized to what we termed as coarsened confounding, for which we propose two new algorithms. We develop asymptotic results for the average causal effect estimator as well as providing conditions for consistency. In addition, we provide an asymptotic justification for the variance formulae in Iacus et al. (2011). A bias correction technique is proposed, and we apply the proposed methodology to data from two well-known observational studies.
@@ 1 Introduction
In this article, we develop a new theoretical framework for coarsened exact matching, which we generalize to what we term as coarsened confounding. Unlike what is presented in Iacus et al. (2011), our approach is based on large-sample theory and thus takes a superpopulation point of view in the terminology of Imbens and Rubin (2015). Results from martingale theory (Fleming and Harrington, 2013; Abadie and Imbens, 2012) and empirical process theory (Van Der Vaart and Wellner, 1996) can be used to study the asymptotic behavior of coarsened confounding-based estimators. Thus, this paper develops a formal theoretical justification for the variance estimates for causal effects from coarsened exact matching of Iacus et al. (2011). In addition, we generalize the structure of the causal effect estimator from Iacus et al. (2011) to develop a data-driven partition-based causal effect estimator. This will allow for two new algorithms for causal effect estimation using machine learning. The first will represent an adaptation of the k-means algorithm (Macqueen, 1967) for estimating causal effects. The second will be based on random forests (Breiman, 2001). While the phenomena we observe are similar to those found in Abadie and Imbens (2006) for nearest neighbor matching estimators, the techniques used for analysis and bias correction are different than what is used in that paper. The structure of this paper is as follows. In Section 2, we set up notation, assumptions and review related literature. Section 3.1 describes the coarsened confounding paradigm and develops some results for asymptotic results for coarsened exact matching as well as the new k-means algorithm for causal effect estimation. Section 4 shows some evaluations using real data examples for the causal effect estimators. Section 5 concludes with some discussion.
@@ 2.3 Related literature
Finally, there is a rich literature on matching methods that this paper adds to. Matching algorithms are a set of procedures that attempt to find for each treated observation in the dataset, the control observation that is ‘closest’ in terms of confounder values. There are many methods available for matching, including nearest-neighbor matching, K:1 matching and optimal matching (Rosenbaum, 1989). Abadie and Imbens (2006) studied the theoretical properties and asymptotics of nearest-neighbor matching procedures. They followed this with work studying the asymptotics of matching on the estimated propensity score (Abadie and Imbens, 2016). A difference between coarsened exact matching with nearest-neighbor matching is that the former does not use any information on $T$ or $Y$ in order to generate the random sets. By contrast, for nearest-neighbor matching, one finds the ‘closest’ control for every treated observation, where the distance is defined based on an appropriate chosen metric for the confounders. Abadie and Imbens (2006) used Euclidean distance assuming the confounders were all continuous. They made the following observations:
@@ 3.1 Proposed Methodology
Recall that coarsened exact matching generates strata ${\cal S}_{1},\ldots,{\cal S}_{J}$ using ${\bf X}_{1},\ldots,{\bf X}_{n}$. We can rewrite the average causal effect $\widehat{\tau}_{CEM}$ from (3.1) as solving the estimating equation $U(\tau)=0$, where
@@
3.2 A quantization-based causal effects estimator
@@
In information theory, a key quantity is the rate distortion function (Berger, 2003). It summarizes how much information can be preserved using data compression methods. It corresponds to the entropy change after data compression. Coarsened exact matching corresponds to taking confounders and recoding it as binary indicators indexing the hypercubes. Assuming $M$ levels for each of the $p$ confounders, the data compression, is approximately $p\log M$. By contrast, $k-$means takes the $n$ observations and maps them to $k$ cluster means, which suggests that the data compression is effectively $\log k$. Note that the data compression for $k-$means clustering is independent of dimension of the confounders. By contrast, the data compression scales linearly in the number of confounders for coarsened exact matching.
@@ 3.4 Asymptotic Analysis
Thus, formulae (3.8) and (3.9) allow for both CEM and the k-means estimator with the attendant standard errors. More generally, we can allow for ${\cal S}_{j}$ $(j=1,\ldots,J)$ to be based on any data-driven algorithm for partitioning based on ${\bf X}$. However, it cannot be based on $T$ or ${\bf Y}$. Having done this, the next question is to understand the asymptotic properties of the estimators. We note the following decomposition for $\widehat{\tau}_{S}-\tau$:
@@
Thus, Theorem 2 provides a formal justification of the variance estimators proposed by Iacus et al. (2011) for coarsened exact matching. Note that our theorem is more general and shows the asymptotic normality for the various data-partitioned causal effect estimators we have proposed: coarsened exact matching, k-means clustering and random forests clustering. In Theorem 2, we also provide a sufficient condition for the bias term in (3.10) to be asymptotically negligible. We note in passing that an alternative to using empirical process theory results to characterize the nature of the bias term in (3.10) would be to use sample splitting/cross-fitting (Zivich and Breskin, 2021) to estimate the bias term directly. Many authors (e.g., Chernozhukov et al. (2018)) have shown that based on sample spltting, one can weaken the smoothness conditions necessary for the bias term to converge to zero in probability. We do not pursue that here.
@@
We now seek to mention some other implications of the martingale theory results we have presented here. First, the construction of the filtrations and $\sigma-$algebras in the theory mean that rules for creating strata that are based on ${\bf X}$ only will have the same theoretical properties as those demonstrated here. We described propensity score subclassification as a related methodology in § 2.3. Our theory would apply to subclassification estimators with a known propensity score. In this case, strata construction would proceed based on $e({\bf X})=P(T=1|{\bf X})$, which is simply a function of ${\bf X}$. However, if we were to fit a model for the propensity score to data, then constructing strata based on $\widehat{e}({\bf X})$ would not fall under our framework. While the notion of using population or known propensity scores do not seem reasonable in practice, we note in passing that much of the work in the seminal paper of Rosenbaum and Rubin (1983) effectively works with $e({\bf X})$ and not $\widehat{e}({\bf X})$. Finally, we do wish to point out that this approach to causal effect estimation avoids propensity scores, but in effect does not model $T|{\bf X}$ directly. Thus, it represents a very different theoretical framework than most causal effect-based estimators, which use estimating functions and semiparametric theory results (Tsiatis, 2006). However, we feel there is merit in understanding the theoretical basis for coarsened confounding, and we believe there are extensions of the approach that could handle more complicated confounding structures, which we leave to future investigations.
@@ 3.6 Bias correction
We instead propose a novel bias-correction approach that is inspired by the simulation-extrapolation approach from the measurement error modeling literature (Carroll et al., 1996), which is commonly abbreviated as SIMEX. In measurement error model, the SIMEX approach fits a sequence of parameter estimates with a given measurement error and extrapolate a parameter estimate based on assuming the measurement error variance approaches zero. For our setting, we will use $J$, the total number of strata, in an analogous way to the measurement error variance. For a given value of $J$, the total number of strata, we compute the causal effect estimators in §3. We then vary $J$ and plot the estimates as a function of $J^{-1}$. Given that we expect consistency in the limit as $J\rightarrow\infty$, we would fit a linear regression of the effect estimates as a function of $J^{-1}$. We would then extrapolate a causal effect estimate at zero. We would employ the same extrapolation procedure with the variance of the estimated causal effect.
@@ 4.1 Right-heart catheterization study
There are 50 confounders in the dataset. When we attempted to match on all confounders using the method in  Iacus et al. (2011), we were unable to get the algorithm to run or to create matched strata with treatment and control observations. Next, we used the K-means based quantization to construct strata. A sequence of K-means results with varying values of $K$ were fit, combined with the bias correction method described in § 3.6. This yielded an average causal effect estimate of -0.55 with a standard error of 0.29. Given that the outcome is binary (1 for survival past 30 days, 0 for death within 30 days), our method reveals that RHC is associated with a causal risk difference of decreasing 30-day survival by 0.55; this effect is significant.
We also applied the random forests-based approach for creating strata, as it allows for more distributional robustness in the confounders. Based on the computed proximity distance matrix, we constructed confounders using hierarchical clustering varying the location where the dendrogram is cut. We again apply the extrapolation method to get a causal risk difference of -0.63 with a standard error of 0.34. Based on the Z-statistic, the random forest approach gives almost identical answers to the k-means analysis.
@@ 4.2 Cesarean observational study
We begin by applying coarsened exact matching. The algorithm yielded a set of 45 strata, with one stratum containing only treated observations. This yielded a risk difference estimate of -1.80, with an attendant standard error of 1.25. Next, we applied the proposed k-means clustering estimator for the risk difference. This yields an causal risk difference of -0.04 with a standard error of 0.52. Finally, the random forests-based causal effect estimate is -0.21 with a standard error of 0.14. Thus, while we find variation in the parameter estimates across the three methods, the direction of the effect estimate is the same. In addition, all three methods do not show evidence for statistical significance for $\alpha=0.05$.
@@ 5 Discussion
The coarsened confounding method discussed in this article could provide consistent estimators for the average causal effect with an asymptotic justification of the number of strata $J\equiv J_{n}\rightarrow\infty$ as the sample size tends to infinity. As has been noted by certain authors (Black et al., 2020), there is a bias in applying coarsened exact matching for real analyses with finite-sample datasets. In order to address this, we introduced a novel bias correction process inspired by the SIMEX (Carroll et al., 1996). In brief, our coarsened confounding method and bias correction process is a comprehensive and effective workflow for approximately consistent casual effect estimation. Future work will seek to provide asymptotic justification for our proposed methodology.
@@
However, there are some issues that still need addressing for real analysis. The first limitation involves how to choose the right variables to represent the population in matching. The effectiveness of matching usually relies on the number of observations retained after matching. The number of variables selected to represent the cohort should be evaluated since the number of matching will reduce as the number of variables increases.
@@
The logic of the binning strategy is also critical here. How to find the balance between the similarity of cohorts and the generalizability in population is worth exploring further. The balance between the number of strata and the efficiency of matching estimation is a prior for the coarsened confounding. In our work, we suggest a consistency properity with the number of strata and sample size both tending to infinity. It is not directly obvious what to do in finite samples. As an alternative to the CEM approach to finding strata, we introduced K-means algorithm for confounder strata clustering, as well as the random forests algorithm. However, these two algorithms do not immediately output a measure of balance. In some articles, an $L_{1}$ statistic was developed as an indicator for balance (Iacus et al., 2009). This statistic ranges from zero to one, where zero refer to the perfect balance with equal proportion of treatment and cohort observations in each stratum, while one refers to the mutually exclusive situation in each stratum. Based on our approach of bias correction process with simulation, we can simulate the range of $L_{1}$ values from zero to one and evaluate it with the different number of strata. We leave this to future work. We could also use this approach with another statistic named Least Common Support (LCS) which indicates the percentage of strata based on different number of observations (Iacus et al., 2009).
@@
A reinterpretation of coarsened exact matching, described in §3.2, is taken in the paper. It exploits an encoding paradigm in which confounders are converted into code vectors, which index observations from which causal effects are computed. Recently, a full encoding-decoding paradigm for causal inference was described in Liu et al. (2024). This leads into deeper connections between causal inference, information theory and machine learning that we intend to explore in future work.

### v2

# Generalized coarsened confounding for causal effects: a large-sample framework

@@ Abstract
There has been widespread use of causal inference methods for the rigorous analysis of observational studies and to identify policy evaluations. In this article, we consider a class of generalized coarsened procedures for confounding. At a high level, these procedures can be viewed as performing a clustering of confounding variables, followed by treatment effect and attendant variance estimation using the confounder strata. In addition, we propose two new algorithms for generalized coarsened confounding. While  Iacus et al. (2011) developed some statistical properties for one special case in our class of procedures, we instead develop a general asymptotic framework. We provide asymptotic results for the average causal effect estimator as well as providing conditions for consistency. In addition, we provide an asymptotic justification for the variance formulae in Iacus et al. (2011). A bias correction technique is proposed, and we apply the proposed methodology to data from two well-known observational studies.
@@ 1 Introduction
In many practical settings, we will have large numbers of confounding variables, say 50 or more. For these settings, coarsened exact matching will suffer from some issues. First, there is a chance that the common support assumption between treatment and control groups will be violated. Second, the coarsened exact matching algorithm will potentially overcoarsen, leading to strata in which there are no observations, or observations from only one group. In the latter scenario, the resulting strata are discarded from the calculation of the treatment effect as well as the attendant standard error. Based on our analysis in Section 3.1, we find that the coarsened exact matching causal effect estimator is in fact a data-adaptive strata-based estimator. This means that the confounders are used to built strata, from which responses in the treatment and control groups are compared. This representation also allows us to study the asymptotic properties of the causal effect estimator and its variance more carefully than has been previously done in the literature.
In this article, we develop a new theoretical framework we term generalized coarsened confounding, of which coarsened exact matching is a special case. Unlike what is presented in Iacus et al. (2011), our approach to inference is based on large-sample theory and thus takes a superpopulation point of view in the terminology of Imbens and Rubin (2015). Results from martingale theory (Fleming and Harrington, 2013; Abadie and Imbens, 2012) and empirical process theory (Van Der Vaart and Wellner, 1996) are used to study the asymptotic behavior of generalized coarsened confounding-based estimators. While our results are relatively general, this paper develops a formal theoretical justification for the variance estimates for causal effects from coarsened exact matching of Iacus et al. (2011). In addition, we develop two new algorithms for generalized coarsened confounding using machine learning. The first is an adaptation of the k-means algorithm (Macqueen, 1967) for estimating causal effects. The second is based on random forests (Breiman, 2001).
One implication of our analysis in Section  3.1 is that there will be a bias in our estimates in finite samples. This is because the creation of discrete strata will induce an estimation bias. While a similar phenomenon was observed for nearest neighbor matching-based causal effect estimation by Abadie and Imbens (2006), we cannot use their techniques for analysis and bias correction. This necessitates a new approach to bias correction that we discuss in Section 3.6. The structure of this paper is as follows. In Section 2, we set up notation, assumptions and review related literature. Section 3.1 describes the coarsened confounding paradigm and develops some results for asymptotic results for coarsened exact matching as well as the new k-means algorithm for causal effect estimation. Section 4 shows some evaluations using real data examples for the causal effect estimators. Section 5 concludes with some discussion.
@@ 2.3 Related literature
Finally, there is a rich literature on matching methods that this paper adds to. Matching algorithms are a set of procedures that attempt to find for each treated observation in the dataset, the control observation that is ‘closest’ in terms of confounder values. There are many methods available for matching, including nearest-neighbor matching, K:1 matching and optimal matching (Rosenbaum, 1989). Once the matched sets are constructed, a variety of approaches can be used to estimate the causal effected in the matched dataset. A nice summary of methods for estimating causal effects in matched datasets can be found in Section 5 of Stuart (2010). Proposed solutions include regression adjustments, weighted analysis approaches and hybrid combinations thereof.
Abadie and Imbens (2006) studied the theoretical properties and asymptotics of nearest-neighbor matching procedures. They followed this with work studying the asymptotics of matching on the estimated propensity score (Abadie and Imbens, 2016). A difference between coarsened exact matching with nearest-neighbor matching is that the former does not use any information on $T$ or $Y$ in order to generate the random sets. By contrast, for nearest-neighbor matching, one finds the ‘closest’ control for every treated observation, where the distance is defined based on an appropriate chosen metric for the confounders. Abadie and Imbens (2006) used Euclidean distance assuming the confounders were all continuous. They made the following observations:
@@ 3.1 Review and Structure Identification
The reinterpretation that we will heavily leverage in this article is that coarsened exact matching generates strata ${\cal S}_{1},\ldots,{\cal S}_{J}$ using ${\bf X}_{1},\ldots,{\bf X}_{n}$. We can rewrite the average causal effect $\widehat{\tau}_{CEM}$ from (3.1) as solving the estimating equation $U(\tau)=0$, where
@@
We note the centrality of equation (3.3) to our generalized coarsened confounding framework. Provided we have an algorithm ${\cal M}$ that takes as its input ${\bf X}_{1},\ldots,{\bf X}_{n}$ and provides as output the strata ${\cal S}_{1},\ldots,{\cal S}_{J}$, we can use (3.3) to compute the average causal effect. While Iacus et al. (2009) use hypercubes based on the components of ${\bf X}$ to define strata, in practice any algorithm for clustering confounders could be used. We will study two other algorithms in this paper, k-means and random forests clustering. What also is evident from (3.3) is that the asymptotic behavior of the estimator will necessitate studying the asymptotic behavior of the sets ${\cal S}\equiv({\cal S}_{1},\ldots,{\cal S}_{J})$, and this is what we do in Section 3.4. We next describe two machine-learning based algorithms for clustering.
3.2 A quantization-based causal effects estimator using k-means
@@
In information theory, a key quantity is the rate distortion function (Berger, 2003). It summarizes how much information can be preserved using data compression methods. It corresponds to the entropy change after data compression. Coarsened exact matching corresponds to taking confounders and recoding it as binary indicators indexing the hypercubes. Assuming $M$ levels for each of the $p$ confounders, the data compression, is approximately $p\log M$. By contrast, $k-$means takes the $n$ observations and maps them to $k$ cluster means, which suggests that the data compression is effectively $\log k$. Note that the data compression for $k-$means clustering is independent of dimension of the confounders. By contrast, the data compression scales linearly in the number of confounders for coarsened exact matching. In information theory, a major goal is to devise procedures that maximize information compression while at the same time preserving information. With respect to the former, our k-means algorithm offers an advantage relative to coarsened exact matching.
@@ 3.4 Asymptotic Analysis
Thus, formulae (3.8) and (3.9) allow for CEM, the k-means estimator and random forests-based clustering with the attendant standard errors. More generally, we can allow for ${\cal S}_{j}$ $(j=1,\ldots,J)$ to be based on any data-driven algorithm for partitioning based on ${\bf X}$. However, it cannot be based on $T$ or ${\bf Y}$. Having done this, the next question is to understand the asymptotic properties of the estimators. We note the following decomposition for $\widehat{\tau}_{S}-\tau$:
@@
Remark 1: In Theorem 2, we also provide a sufficient condition for the bias term in (3.10) to be asymptotically negligible. We note that condition (3.13) in Theorem 2 is in fact an asymptotic condition on the complexity of the sets that are induced by the partitioning algorithm. We now discuss the implications of the condition for coarsened exact matching, k-means and random forests based clustering. For coarsened exact matching, we assume that for each of the $p$ covariates, we partition them into $M$ bins. Some standard combinatorial arguments (e.g., p. 367 of  Devroye et al. (2013)) can be used to show that for this set of partitions, ${\cal G}_{n}^{(M)}$,
$\Delta_{n}({\cal G}_{n}^{(M)})\leq 2^{M^{p}}\binom{n+M}{n}^{p}.$
We thus have that (3.13) is satisfied when
$\lim_{n\rightarrow\infty}\frac{M_{n}^{p}}{n}=0.$
For the $k-$means partition approach to clustering, we leverage the work of Lugosi and Nobel (1996), who demonstrate concentration bounds for Voronoi partitions of points in $R^{p}$. Based on our notation and their results, defining $J_{kn}$ to be the closest neighbors of the $k$th cluster center and allowing for dependence on sample size, if
$J_{kn}\rightarrow\infty\ \ \text{and}\ \ \frac{J^{2}_{kn}\log n}{n}\rightarrow 0$
as $n\rightarrow\infty$, then (3.13) is satisfied. Finally, for the the random forests approach, we begin by assuming a single tree with at most $S_{n}$ splits, where
$\frac{S_{n}\log n}{n}\rightarrow 0.$
as sample size approaches infinity. Then the results of Lugosi and Nobel (1996) can again be used to verify that (3.13) holds. Random forests now aggregates over $B$ trees, so the condition on $S_{n}$ will again imply that for the aggregated tree, (3.13) holds. We wish to reiterate that condition (3.13) holding is theoretical and relies on asymptotic considerations in which sample size and other tuning parameters are approaching infinity with potential constraints on their interactions.
@@
Theorem 2 provides a formal justification of the variance estimators proposed by Iacus et al. (2011) for coarsened exact matching. Note that our theorem is much more general and shows the asymptotic normality for the various data-partitioned causal effect estimators we have proposed: coarsened exact matching, k-means clustering, and random forests clustering. We note in passing that an alternative to using empirical process theory results to characterize the nature of the bias term in (3.10) would be to use sample splitting/cross-fitting (Zivich and Breskin, 2021) to estimate the bias term directly. Many authors (e.g., Chernozhukov et al., 2018) have shown that based on sample splitting, one can weaken the smoothness conditions necessary for the bias term to converge to zero in probability. We do not pursue that here.
@@
We now seek to mention some other implications of the martingale theory results we have presented here. First, the construction of the filtrations and $\sigma-$algebras in the theory means that rules for creating strata that are based on ${\bf X}$ only will have the same theoretical properties as those demonstrated here. We described propensity score subclassification as a related methodology in § 2.3. Our theory would apply to subclassification estimators with a known propensity score. In this case, strata construction would proceed based on $e({\bf X})=P(T=1|{\bf X})$, which is simply a function of ${\bf X}$. However, if we were to fit a model for the propensity score to data, then constructing strata based on $\widehat{e}({\bf X})$ would not fall under our framework. While the notion of using population or known propensity scores do not seem reasonable in practice, we note in passing that much of the work in the seminal paper of Rosenbaum and Rubin (1983) effectively works with $e({\bf X})$ and not $\widehat{e}({\bf X})$. Finally, we do wish to point out that this approach to causal effect estimation avoids propensity scores, but in effect does not model $T|{\bf X}$ directly. Thus, it represents a very different theoretical framework than most causal effect-based estimators, which use estimating functions and semiparametric theory results (Tsiatis, 2006). However, we feel there is merit in understanding the theoretical basis for generalized coarsened confounding, and we believe there are extensions of the approach that could handle more complicated confounding structures, which we leave to future investigations.
@@ 3.6 Bias correction
We instead propose a novel bias-correction approach that is inspired by the simulation extrapolation approach from the measurement error modeling literature (Carroll et al., 1996), which is commonly abbreviated as SIMEX. In measurement error modeling, the SIMEX approach fits a sequence of parameter estimates with a given measurement error and extrapolates a parameter estimate based on assuming the measurement error variance approaches zero. For our setting, we will use $J$, the total number of strata, in an analogous way to the measurement error variance. Here is the outline for the approach.
  1. 1.
    Create a grid of values of $J$.
  2. 2.
    For each value of $J$ on the grid, compute the estimated average causal effect and the variance.
  3. 3.
    Fit a linear regression of the estimates as a function of $J^{-1}$.
  4. 4.
    Use as the proposed estimate the predicted value at $J^{-1}$. If the previous step was fit using the lm function in R and and saved as an object foo, then one would use predict(foo,invJ = 0) to obtain the estimate.
The idea behind the proposed bias correction approach is that we expect unbiased causal effect estimation to occur when $J$, the number of strata, approaches infinity, or equivalently when $J^{-1}$ approaches zero. For a given value of $J$, our theory from section 3 yields that the limiting distribution is asymptotically normal. Thus, we are making the assumption that the interpolated value will also have an asymptotic normal distribution. A heuristic argument is that since the causal effect estimator in step 5 is a linear combination of the estimates at each value of $J$, it should also be asymptotically normal. Formal justification of this result is beyond the current scope of the paper and is currently under investigation.
@@ 4 Examples
For all the examples, we assume that for Wald statistics, we can apply the methods from Sections 3.4 and 3.6 so that a normal distribution under the null hypothesis holds. In addition, we employ a significance level of 0.05.
@@ 4.1 Right-heart catheterization study
There are 50 confounders in the dataset. When we attempted to match on all confounders using the method in  Iacus et al. (2011), we were unable to get the algorithm to run or to create matched strata with treatment and control observations. Next, we used the K-means based quantization to construct strata. A sequence of K-means results with varying values of $K$ were fit, combined with the bias correction method described in § 3.6. This yielded an average causal effect estimate of -0.041 with a standard error of 0.01. Given that the outcome is binary (1 for survival past 30 days, 0 for death within 30 days), our method reveals that RHC is associated with a causal risk difference of decreasing 30-day survival by 0.04; this effect is significant $(p=0.003)$.
We also applied the random forests-based approach for creating strata, as it allows for more distributional robustness in the confounders. Based on the computed proximity distance matrix, we constructed confounders using hierarchical clustering varying the location where the dendrogram is cut. We again apply the extrapolation method to get a causal risk difference of -0.021 with a standard error of 0.014. Based on the Z-statistic, the random forest approach gives less significant answers to the k-means analysis, with a p-value of $0.13$. However, the directionality of the causal effect from random forests aligns with that of k-means. We have included the R code for the analysis in the Appendix.
@@ 4.2 Cesarean observational study
We begin by applying coarsened exact matching. The algorithm yielded a set of 45 strata, with one stratum containing only treated observations. This yielded a risk difference estimate of $8.62\times 10^{-5}$, with an attendant standard error of $0.005$ (p-value = $0.98$). Next, we applied the proposed k-means clustering estimator for the risk difference. This yields a causal risk difference of $0.004$ with a standard error of $0.005$ (p-value = $0.44$) . Finally, the random forests-based causal effect estimate is $5.27\times 10^{-4}$ with a standard error of $0.005$ (p-value = $0.92$). We note that the magnitudes of the k-means and random forests estimator are substantially larger in magnitude relative to the original CEM estimate of the average causal effect. We note that while the former two are bias corrected, the CEM estimate is not. We also ran a separate analysis with k-means clustering using $K=45$ to match the number of strata in CEM, followed by the two-sample difference in means to match CEM. This yielded an estimate of $2.2\times 10^{-4}$ with a standard error of $0.005$. Reasons for the discrepancy between the CEM and the other approaches include the lack of bias correction for CEM as well as variation in the grouping of observations into strata. Thus, while we find variation in the parameter estimates across the three methods, the direction of the effect estimate is the same. In addition, all three methods do not show evidence for statistical significance for $\alpha=0.05$. We have included the code and results for these results in the Appendix.
@@ 5 Discussion
The generalized coarsened confounding method discussed in this article can provide consistent estimators for the average causal effect with an asymptotic justification of the number of strata $J\equiv J_{n}\rightarrow\infty$ as the sample size tends to infinity. As has been noted by certain authors (Black et al., 2020), there is a bias in applying coarsened exact matching for real analyses with finite-sample datasets. We demonstrated this formally in Section 3. In order to address the induced bias, we have introduced a novel bias correction process inspired by the SIMEX procedure from Carroll et al. (1996). In summary, our generalized coarsened confounding methodology and bias correction process provides a workflow for approximately consistent casual effect estimation. Future work will seek to provide asymptotic justification for the bias correction methodology.
@@
However, there are some issues that still need to be addressed in real-data analyses. The first involves how to choose the right variables to represent the population in matching. The effectiveness of matching usually relies on the number of observations retained after matching. The number of variables selected to represent the cohort should be evaluated since the number of matched observations will decrease as the number of matching variables increases.
@@
The choice of binning strategy is also critical here. How to find the balance between the similarity of cohorts and the generalizability in population is worth exploring further. In our work, we suggest a consistency properity with the number of strata and sample size both tending to infinity. It is not directly obvious what to do in finite samples. As an alternative to the CEM approach to finding strata, we introduced K-means algorithm for confounder strata clustering, as well as the random forests algorithm. However, these two algorithms do not immediately output a measure of balance. In some articles, an $L_{1}$ statistic was developed as an indicator for balance (Iacus et al., 2009). This statistic ranges from zero to one, where zero refers to perfect balance with equal proportion of treatment and control observations in each stratum, while one refers to the mutually exclusive situation in each stratum. Based on our bias correction approach, we can simulate the range of $L_{1}$ values from zero to one and evaluate it with the different number of strata. We leave this to future work. We could also use this approach with another statistic named Least Common Support (LCS) which indicates the percentage of strata based on different number of observations (Iacus et al., 2009).
@@
A reinterpretation of coarsened exact matching, described in §3.2, is used in the paper. It exploits an encoding paradigm in which confounders are converted into code vectors, which index observations from which causal effects can be computed. Recently, a full encoding-decoding paradigm for causal inference was described in Liu et al. (2024). This leads into deeper connections between causal inference, information theory and machine learning that we intend to explore in future work.